OpenStudy (anonymous):

Check my answer? Simplified form of (x-2) (2x+3). I got 2x^2 - x - 6.

OpenStudy (anonymous):

2x*x=2x^2 3*x=3x -2*2x=-4x 3*-2=-6 I get 2x^2-x-6 when simplified. So correct.

OpenStudy (anonymous):

Yay :D Thanks, I wasn't sure.

OpenStudy (anonymous):

@curiouscrab could you check this answer for me? Simplify (3x+2) (4x-3) I got 12x^2+x-6

OpenStudy (anonymous):

Your answer is correct

OpenStudy (anonymous):

Thank you @sensuelle1985 :D

OpenStudy (anonymous):

@sensuelle1985 , could you check 2 more for me?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh so just forget me right @Destinymasha hey baby

OpenStudy (destinymasha):

ermm........

OpenStudy (anonymous):

@Destinymasha nah nah its a joke dude

OpenStudy (anonymous):

Simplify (4p-2) (p-4) I got 4p^2-18p+8

OpenStudy (anonymous):

lol you're famous @Destinymasha on this site atleast

OpenStudy (anonymous):

I'm working on the last one now.

OpenStudy (anonymous):

@SunSeto your answer is correct

OpenStudy (anonymous):

K give me a moment while I work on the last one.

OpenStudy (anonymous):

not a problem

OpenStudy (anonymous):

what about me

OpenStudy (anonymous):

no love?

OpenStudy (anonymous):

I just wanted him to check my answers lol, I'm not asking for them.

OpenStudy (anonymous):

Okay @sensuelle1985 , I got the answer but I need it simplified. I'm not sure how to go about it though. I got 6x^3+2x^2+12x+4x+9x-2

OpenStudy (anonymous):

I mean -3

OpenStudy (anonymous):

what is the equation

OpenStudy (anonymous):

The equation is (2x^2+4x-3) (3x+1)

OpenStudy (anonymous):

answer will be \[6x^3+14x^2 -5x-3\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

kendrick lamar ftw...........sorry guys im only here to troll

OpenStudy (anonymous):

lolololol

OpenStudy (anonymous):

u ad

OpenStudy (anonymous):

u mad

OpenStudy (anonymous):

bumbump

OpenStudy (anonymous):

Thank you very much @sensuelle1985 , I've just never multiplied a trinomial and a binomial before.

OpenStudy (anonymous):

ha sunseto what a funny name

OpenStudy (anonymous):

@SunSeto @SunSeto

OpenStudy (anonymous):

@SunSeto anytime, that's why we are here for.

OpenStudy (anonymous):

@sensuelle1985

OpenStudy (anonymous):

yes @yonkou

OpenStudy (anonymous):

@sensuelle1985 what up